### v1

# Learning Task-adaptive Quasi-stiffness Control for a Powered Transfemoral Prosthesis

Abstract
While significant advancements have been made in the mechanical and task-specific controller designs of powered transfemoral prostheses, developing a task-adaptive control framework that generalizes across various locomotion modes and terrain conditions remains an open problem. This study proposes a task-adaptive learning quasi-stiffness control framework for powered prostheses that generalizes across tasks, including the torque-angle relationship reconstruction part and the quasi-stiffness controller design part. Quasi-stiffness is defined as the slope of the human joint’s torque-angle relationship. To accurately obtain the torque-angle relationship in a new task, a Gaussian Process Regression (GPR) model is introduced to predict the target features of the human joint’s angle and torque in the task. Then a Kernelized Movement Primitives (KMP) is employed to reconstruct the torque-angle relationship of a new task from multiple human demonstrations and estimated target features. Based on the torque-angle relationship of the new task, a quasi-stiffness control approach is designed for a powered prosthesis. Finally, the proposed framework is validated through practical examples, including varying speed and incline walking tasks. The proposed framework has the potential to expand to variable walking tasks in daily life for the transfemoral amputees.
I Introduction
Transfemoral amputees suffer ambulation difficulties when performing demanding locomotion modes, such as climbing ramps and stairs. Powered transfemoral prostheses have the potential to restore the lost function of the joints and enhance the mobility of the amputees [1, 2]. While significant progress has been made in both the mechanical designs [3, 4, 5, 6, 7, 8, 9, 10] and controller designs [11, 12, 13, 14, 15, 16, 17] of powered prostheses over the past decades, developing a prosthesis control framework that generalizes across a variety of locomotion modes and terrain conditions remains a challenge [2].
Impedance-based control is the most commonly used approach in powered prostheses because of its ability to reproduce natural walking. These methods segment the gait cycle into several sub-phases, each performing an impedance control law. Previous researchers manually tuned the impedance parameters for each task, such as flat walking [3], ramp climbing [18], and stair climbing [19]. However, the heuristic parameter tuning procedure for a specific task takes at least several hours [20]. As shown in Fig. 1, transfemoral amputees may perform different tasks (locomotion modes and terrain conditions) in daily walking. The time required for tuning a powered prosthesis for each task and each patient presents a major barrier to its clinical feasibility [21]. To reduce the tuning time, reinforcement learning (RL) is implemented in [14, 22] to automatically tune the impedance parameters. Nevertheless, these approaches are limited to flat walking task and knee joint only at present. Recent studies have achieved prostheses handling variable-task walking. Data-driven, impedance-based controllers are respectively designed for varying speeds and inclines walking [20], stair climbing [23], and sit/stand [24]. An adaptive stair controller enables transfemoral amputee patients to climb stairs of varying stair heights [25]. These controllers are still required to extend to a unified framework for variable-task walking. To date, developing powered transfemoral prostheses controllers that generalize across tasks remains an open problem.
Quasi-stiffness control is a subset of the impedance-based control strategies[2]. This strategy determines the impedance parameters of each sub-phase from the torque-angle relationship of human’s lower limb joints. The slope of the torque-angle relationship is called ”quasi-stiffness” [2, 26]. The desired prosthesis torque is then determined by the quasi-stiffness and equilibrium angle as . The torque–angle relationships can be modified for quasi-stiffness control of powered prostheses in different tasks. For example, the stair controller modulates the torque-angle relationship based on the stair height for stair climbing tasks. In addition, it is reported that quasi-stiffness control strategies have the potential to improve gait symmetry and achieve speed-adaption without subject-specific or speed-specific tuning [21]. Inspired by the quasi-stiffness control strategy, this study proposes a learning task-adaptive quasi-stiffness control framework that generalizes across a variety of tasks. Two issues are presented and need to be resolved. The first one is to automatically construct joints’ torque-angle relationship in a new task. The second one is to design quasi-stiffness controllers based on the generated torque-angle relationship.
The torque-angle relationship is determined by the joint angle and torque and is widely used in the design and control of powered exoskeleton [27, 28] and prosthesis [21, 25]. The peak values (named as target features in this study) in the profiles of joint kinematics and kinetics vary with tasks and are often used to represent the joint kinematics and kinetics during walking tasks [14, 22]. However, most torque-angle relationships and target features are obtained in the laboratory by a motion capture system. It is impossible to collect joint kinematics and kinetics for all tasks. This study proposed an imitation learning-based method that can reconstruct the torque-angle relationships of new tasks by learning from multiple demonstrated torque-angle relationships. For example, reconstructing a relationship at a new walking speed from learning the existing demonstrated relationship at different speeds (or a new relationship at a new ramp inclination, etc.). To reconstruct the torque-angle relationship of a new task, the Gaussian Process Regression (GPR) [29, 30] is introduced to predict the target features of the joint angles and torques based on the demonstrated features and task variables (such as speed and terrain parameters). Furthermore, the new task torque-angle relationship can be reconstructed by the target features and human-demonstrated torque-angle relationships. While many fitting methods can use these target features for torque-angle relationship reconstruction, such as polynomial interpolation [31] and Fourier Series fitting [32], these methods consider only trajectory fitting instead of the multiple demonstrated relationships. Instead, learning from multiple demonstrations methods such as probabilistic trajectory methods can be introduced. Gaussian Mixture Model/Regression (GMM-GMR) [33, 34] was applied to reproduce joint angles and learning trajectory distribution through learning from the demonstrated database. However, this method makes it difficult to modify the learned trajectories to fulfill additional new constraints, such as passing through specific target features. A Kernelized Movement Primitives (KMP) [35] method was presented that allows the reproduced learning from demonstration trajectories passing through the specific via-points. Inspired by this, KMP is employed to construct the new torque-angle relationship from multiple demonstrated relationships and target features. After obtaining the torque-angle relationship, the quasi-stuffiness control strategy is introduced to design a task-adaptive controller that generalizes across tasks.
The main contributions are summarized as follows:
1.
A learning-based torque-angle relationship construction method is proposed for various tasks.
2.
A quasi-stiffness control strategy based on the learned torque-angle relationship is presented for powered transfemoral prostheses.
3.
Examples of varying speed and varying incline walking using the proposed quasi-stiffness control framework are conducted and validated through walking experiments with a transfemoral amputee.
II Method
As shown in Fig. 2, the proposed task-adaptive learning quasi-stiffness control framework consists of a learning from human demonstrations part and a quasi-stiffness control part. The learning-based part is composed of two models: a GPR-based target features prediction model and a KMP-based task-specific torque-angle relationship reconstruction model.
II-A Task Estimation
Our previous studies and some other studies presented various task estimation methods, such as vision-based methods [36, 37] for terrain condition estimation and IMU (inertial measurement unit)-based methods [20, 38, 39] for speed estimation. These methods can be applied to estimate the task parameter (where denotes the walking speed, represents the terrain parameters) for the proposed framework.
II-B GPR-based Target Features Prediction
An open-source dataset [40] is implemented as the human demonstrations, which contains tasks including multiple locomotion modes (flat walking, ramp climbing, and stair climbing) and terrain conditions (walking speeds, ramp inclinations, and stair heights). To obtain the new task-specific torque-angle relationship, this study introduces GPR-based learning from the demonstration method for target features prediction, as shown in Fig. 2(b). Considering the limited human demonstration size, GPR, a nonparametric kernel-based probabilistic model, adeptly balances fitting accuracy with model complexity, mitigating the risk of overfitting in small-size training samples. Assume there are demonstrations and define the train data as
where denotes the number of the features; is the task parameter of the -th demonstration (-th task). is the output features (joint angles/torques) of the -th demonstration. For the task , the GPR [29, 30] is intended to transform the input task vector into a output value by , where represents the Gaussian noise satisfying . The training dataset can be described by the GPR model as
where represents the mean function and is set to zero [30]; denotes the covariance function (also called Gaussian kernel) and can be selected in [29]. The target features’ prior distribution in satisfies
where is the covariance matrix. To generate prediction for a new task input , the predicted function is following a joint Gaussian distribution:
The posterior distribution of the prediction at a new input is given by
The predicted mean value and corresponding variance of the conditional distribution can be given by [30]
where , denotes the covariance matrix between training and new input, and denotes the covariance between the new input vector.
Thus, the new output is predicted as the mean values. By employing the GPR method, the new task’s target features including joint angles and joint torques can be represented as where represents the number of target features.
II-C KMP-based Torque-angle relationship Reconstruction
After obtaining the kinematics and kinetics target features of the new task, the KMP is employed to reconstruct the torque-angle relationship with the target features of the new task, as shown in Fig. 2(c). Denote the set of demonstrated torque-anlge relationship by where represents the joint kinematics and is the joint kinetics. and denote the length of the torque-angle relationship and the number of the demonstrations. GMM [34] is introduced to encode the demonstrated dataset and estimate the relationship probability distribution
where , , and represent prior probability, mean, and covariance, respectively. is the number of the Gaussian component. Furthermore, a probabilistic reference relationship can be retrieved by GMR
where and represent the reference torque and angle respectively, and denote mean and variance respectively. As mentioned before, we expect that the new torque-angle relationship can pass through the target features of the new task. Therefore, KMP is employed to learn from demonstrations and fulfill the additional constraints. Considering a parametric torque-angle relationship can be formulated by KMP as
where is a -dimensional basis function and is defined in [35]. is the weight vector and can be assumed to be normally distributed following . Thus, the probability distribution of the parametric torque-angle relationship stratifies
The new torque-angle relationship is expected to pass through the target features. The target features is assumed satisfying and
Finally, KMP requires matching the reference torque-angle relationship with the parametric torque-angle relationship and the target features. Inspired by [35, 41, 42], to obtain and , the Kullback-Leibler (KL) divergence is introduced to design the cost function as
where denotes the KL divergence between the two probability distributions. A detailed solution to minimization of the cost function can be found in [35]. By employing the KMP, the torque-angle relationship of the new task can be reconstructed and applied in the quasi-stiffness control.
II-D Quasi-stiffness Estimation and Controller Design
For the powered prostheses, the desired torque determined by the ”quasi-stiffness” control law is defined as
where denotes the quasi-stiffness, represents the equilibrium angle. It is noted that the torque-angle relationship during the stance phase is approximately linear [26, 43]. Inspired by this, as shown in Fig. 2(d), this study applies a linear regression to estimate the quasi-stiffness of the torque-angle relationship, and the equilibrium angle is equal to the value of the intersection point between the regression equation and the x-axis. For the prosthesis control, we divide the gait phase into four sub-phases: stance flexion, stance extension, swing flexion, and swing extension. The sub-phase switching follows a finite-state machine (FSM), as shown in Fig. 3. For the stance phase, the torque-angle relationship is approximately linear, which can be approximated by a linear equation. While the swing phase is much more complex than the stance phase [43], we choose two linear parts that can achieve swing flexion and extension and use linear functions to approximate these two parts.
III Results
III-A Experimental Setup
The powered transfemoral prosthesis used in this study is shown in Fig. 6(a). Both knee and ankle joints are driven by frameless motors (ILM motor kit, RoboDrive, Seefeld, Germany) with a 50:1 harmonic reducer (CSD-25-50-2A-GR, Nagano Prefecture, Japan). The frameless motors are actuated by drives (G-MOLTWIR50/100SEO, Elmo Motion Control, Petah Tikva, Israel). A 6-axis force/torque sensor (Sunrise Instruments, Nanning, China) is employed to measure the (ground reaction forces) GRFs.
A transfemoral amputee patient is invited to participate in the experiments to evaluate the proposed learning quasi-stiffness control framework. The participant provided informed consent, and the Southern University of Science and Technology Institutional Review Board approved the experiments.
III-B Experimental Results of the Proposed Framework
Due to the space limitations, this study presents an example of varying walking speed tasks to show how the learning quasi-stiffness control framework works. The learning process of the GPR-based target feature estimation and KMP-based torque-angle relationship construction is offline on the joint kinematics and kinetics obtained from the open-source dataset [40].
1) Target Features Estimation: Fig. 2(b) shows the experimental results of target features prediction, where the blue dots denote the demonstrated target features and the purple dots represent the predicted target features of a new task (new speed). The GPR-based target features prediction model is imported into the embedded platform of the powered prosthesis and used for feature prediction of new tasks.
2) Torque-angle relationship Reconstruction: Fig. 4 shows the torque-angle relationship learning and reconstruction in a new task. The distribution of the demonstrated torque-angle relationship can be learned by GMM. Based on the learned GMM, the reference torque-angle relationship can be reproduced by GMR. As shown in Fig. 4(a), the blue torque-angle curves denote the human demonstrations of different speed walking tasks, and the yellow torque-angle curve represents the reference torque-angle relationship reproduced by GMR. As shown in Fig. 4(b), the yellow dashed curve represents a torque-angle relationship reconstructed by the demonstrated torque-angle relationships (blue curves) and target features (purple dots) in a new task.
3) Quasi-stiffness Estimation: Fig. 5 shows using linear functions to approximate several parts of the knee and ankle torque-angle relationships in the varying speed walking tasks. Table I shows the estimated stiffness and equilibrium angles from the torque-angle relationships of different speed tasks and these impedance parameters are used in the powered transfemoral prosthesis experiments.
4) Performance of the Proposed Method in Experiments: Fig. 6(b) shows a series of photos of a transfemoral amputee performing varying speed and inclines walking tasks. Fig. 7 presents that the kinematic and kinetics trajectories produced by the proposed method are similar to the able-bodied references. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which also indicate that the quai-stiffness can improve gait symmetry. Fig. 8 shows the root mean squared error (RMSE) in the powered prosthetic joint trajectories relative to able-bodied trajectories and intact leg trajectories during the walking experiments. The lower RMSE of the observed kinematics with respect to the intact leg kinematics indicates that the proposed quasi-stiffness control achieves better joint kinematic symmetry.
| Task (m/s) | Joint | Quasi-stiffness (Nm/deg) | Equilibrium Angle (deg) | ||||||
| 0.4 | knee | 4.13 | 0.59 | 0.08 | 0.19 | 12 | 22 | 30 | 15 |
| ankle | 0.82 | 6.14 | 0.53 | 0.59 | -5 | -2 | 8 | -8 | |
| 0.6 | knee | 5.68 | 0.70 | 0.17 | 0.26 | 12 | 21 | 57 | 17 |
| ankle | 0.76 | 5.05 | 0.53 | 0.64 | -5 | 0 | 8 | -8 | |
| 0.8 | knee | 5.79 | 0.79 | 0.27 | 0.45 | 13 | 22 | 69 | 18 |
| ankle | 0.78 | 4.23 | 0.45 | 0.69 | -6 | -4 | 9 | -7 | |
| 1.0 | knee | 5.85 | 0.85 | 0.36 | 0.52 | 13 | 22 | 76 | 20 |
| ankle | 1.03 | 3.89 | 0.36 | 0.72 | -5 | -7 | 10 | -8 | |
IV Discussion
IV-A Evaluation of the Proposed Control Method
This work proposed a learning task-adaptive quasi-stiffness controller for a powered transfemoral prosthesis that generalizes across tasks. The proposed method has the potential to quickly and autonomously configure the impedance parameters for variable tasks. Experimental results and walking performance (see the accompanying video) highlight its capacity to replicate able-bodied biomechanics across variable tasks. As shown in Fig. 7, the observed kinematic and kinetic profiles of the powered prosthetic joint vary with walking speed and are similar to able-bodied references, such as appropriately varying peak angles, peak moments, and trends of profiles. The powered prosthetic joint trajectories show strong similarities to the intact joint trajectories, which indicates that the quai-stiffness control has the potential to improve gait symmetry. Knee kinematics show the most differences with respect to able-bodied data, which may be because the torque-angle relationship of the knee joint is more complex with fewer linear parts, which inevitably introduces errors when using linear approximations.
IV-B Limitations and Future Work
A limitation of this study is that the torque-angle relationship of the swing phase tends to be more nonlinear and complex as the task becomes more difficult (higher walking speed or more challenging terrain conditions), which means that using linear regression to estimate quasi-stiffness may introduce larger errors. In future work, we will investigate the use of nonlinear regression to estimate the swing phase quasi-stiffness. Nevertheless, this work focuses on designing a powered transfemoral prosthesis control framework that generalizes across various tasks. Experimental results and walking performance (see the accompanying video) preliminarily verify that the framework has the potential to achieve task-adaptive control. Future studies will invite more participants to verify the performance of the proposed framework in outdoor experiments.
V Conclusion
This study proposed a learning task-adaptive quasi-stiffness control framework for a powered transfemoral prosthesis working over variable tasks. Learning from the human demonstration method is developed for the torque-angle relationship reconstruction in a new task, which consists of a GPR-based target features prediction model and a KMP-based trajectories reconstruction model. A transfemoral amputee participant demonstrated the control framework’s ability to generalize across tasks by varying speeds and inclines walking experiments. The experimental results validate that the proposed control framework has the potential to reproduce biomimetic joint torques and angles for a powered prosthesis during variable-task walking.